# Maximizing Asynchronicity in Event-based Neural Networks

### v1

@@ Abstract
Event cameras deliver visual data with high temporal resolution, low latency, and minimal redundancy, yet their asynchronous, sparse sequential nature challenges standard tensor-based machine learning (ML). While the recent asynchronous-to-synchronous (A2S) paradigm aims to bridge this gap by asynchronously encoding events into learned representations for ML pipelines, existing A2S approaches often sacrifice representation expressivity and generalizability compared to dense, synchronous methods. This paper introduces EVA(EVent Asynchronous representation learning), a novel A2S framework to generate highly expressive and generalizable event-by-event representations. Inspired by the analogy between events and language, EVA uniquely adapts advances from language modeling in linear attention and self-supervised learning for its construction. In demonstration, EVA outperforms prior A2S methods on recognition tasks (DVS128-Gesture and N-Cars), and represents the first A2S framework to successfully master demanding detection tasks, achieving a remarkable 47.7 mAP on the Gen1 dataset. These results underscore EVA’s transformative potential for advancing real-time event-based vision applications.
footnotetext: Emails: hhq23@mails.tsinghua.edu.cn, zubic@ifi.uzh.ch, hwh20@mails.tsinghua.edu.cn, hzp22@mails.tsinghua.edu.cn, sdavide@ifi.uzh.ch, wwh@tsinghua.edu.cn
@@ 1 Introduction
Traditional frame-based cameras synchronously capture dense visual intensity data at fixed rates. In contrast, event cameras asynchronously detect pixel-level intensity changes, resulting in event sequences characterized by high temporal resolution (up to 1 $\mu$s) and low spatial redundancy gallego2020event . Due to their ability to capture rapid movements with minimal latency and reduced bandwidth requirements, event cameras have garnered widespread attention in autonomous driving gehrig2024low , robotics he2024microsaccade , and biomedical analysis he2024neuromorphic . However, the asynchronous and sparse nature of event data inherently complicates its direct and efficient use with standard machine learning (ML) algorithms, which typically demand tensor-based inputs. Current approaches to utilize event data in vision tasks predominantly fall into two categories: converting event sequences into dense, image-like representations aggregated over fixed time windows sironi2018hats ; lagorce2016hots ; gehrig2019end ; innocenti2021temporal ; Zubic_2023_ICCV , or employing spiking neural networks (SNNs) he2024neuromorphic . The former often fails to fully exploit the inherent sparsity and rich temporal information, while the latter relies on specialized hardware and encounters challenges in training high-performance models effectively martin2024alert .
To better leverage the fine-grained temporal feature and inherent sparsity of events while benefiting from advancements in standard ML, the asynchronous-to-synchronous (A2S) paradigm has recently emerged martin2024alert . This paradigm focuses on designing an efficient asynchronous encoder that recurrently integrates events into a tensor-based representation on an event-by-event basis in real time. This representation can then be synchronously sampled at flexible, on-demand rates by downstream ML tasks. However, existing A2S methodologies martin2024alert rely on preliminary models with limited expressivity as a compromise for computational efficiency. Consequently, they achieve suboptimal performance on complex tasks compared to dense, tensor-based approaches gehrig2023recurrent ; gehrig2019end ; Zubic_2024_CVPR . Furthermore, these representations are typically learned in a supervised manner, limiting their generalizability across diverse downstream applications.
@@
Figure 1: Overview of the proposed EVA A2S framework. It features an asynchronous LA-based encoder (top middle) producing event-by-event updated representations for synchronous downstream tasks (right). These representations are learned via a self-supervised module (bottom middle) and are event-by-event updated during inference to be sampled on-demand by downstream tasks.
@@
To design an efficient A2S framework with more expressivity and generalizability, we take inspiration from the analogy between events and language Soydan2024S7SA . On one hand, events and language are both organized in sequences. Events record incremental visual changes, similar to words in a sentence, incrementally contributing to the semantic meaning. This inspires us to treat each event as a discrete token and leverage recent advances in natural language processing (NLP), including linear attention (LA) for efficient sequence modeling katharopoulos2020transformers ; yang2023gated ; peng2023rwkv ; sun2023retentive and auto-regressive prediction radford2018improving for self-supervised learning. On the other hand, crucial distinctions between events and language exist, primarily in information density and spatial locality. A single language token inherently possesses significant semantic meaning, whereas an individual event offers limited information, gaining significance primarily through aggregation within its historical context. This distinction steers our focus towards the aggregated global information rather than isolated event instances. Additionally, events typically signify spatially localized changes, contrasting with language, which lacks inherent spatial attributes. This motivates modeling correlations primarily between spatially adjacent events to reduce complexity.
@@
In this work, we introduce a new A2S framework with a more expressive asynchronous encoder, which achieves event-by-event representation updating, and a self-supervised method for highly generalizable representation learning. Our asynchronous encoder is built on the basis of LA, which enables recurrent inference and parallel training. We propose to use the matrix-value hidden state (MVHS) introduced in LA as output to expand the model’s expressivity on the aggregated global information and propose patch-wise encoding to leverage the locality. For ease of implementation and better performance, our model is built on a high-performance open source LA architecture RWKV-6 peng2024eagle , which implements MVHS and is stable to train. We propose a self-supervised method for generalizable representation learning, which consists of two tasks: multi-representation prediction (MRP) and next-representation prediction (NRP). The former encourages the representation to learn diverse information from multiple handcrafted (converted) representations lagorce2016hots ; sironi2018hats comprehensively, and the latter, motivated by next-token prediction strategies prevalent in NLP radford2018improving , pushes the model beyond simple memorization. Our framework is superior to the previous A2S work martin2024alert on recognition tasks, and it is the first A2S framework, to the best of our knowledge, that successfully masters demanding detection tasks, achieving a remarkable 47.7 mAP on the Gen1 dataset. We name our framework as EVent Asynchronous representation learning (EVA). Our contributions include:
@@
1. An asynchronous LA-based encoder architecture derived from RWKV-6, enabling efficient event-by-event representation updating with improved expressivity.
@@
2. A novel multi-task SSL method that learns generalizable representations across various downstream tasks.
@@
3. We present EVA, a new A2S framework that synergizes the asynchronous encoder with self-supervised representation learning. Compared to the previous A2S work, EVA outperforms on recognition tasks and, for the first time, successfully masters demanding detection tasks with a remarkable 47.7 mAP on the Gen1 dataset.
@@ 2.1 Asynchronous event processing
To fully leverage the sparsity and high temporal resolution of event data within deep learning, several recent efforts have focused on asynchronous event processing. These include point cloud approaches sekikawa2019eventnet ; martin2024alert , convolutional methods messikommer2020event ; cannici2019asynchronous , graph-based models li2021graph ; schaefer2022aegnn ; Zubic_2025_CVPR , attention-based mechanisms kamal2023associative , and LSTM-based solutions cannici2020differentiable ; santambrogio2024farse . Specifically, EventNet sekikawa2019eventnet introduces an event-by-event feature update mechanism based on PointNet qi2017pointnet . However, features are encoded into a vector and therefore limited in expressivity. Convolutional methods have explored deriving local update rules of activations in standard cannici2019asynchronous or sparse messikommer2020event convolution layers. While they can leverage sparsity and reduce redundant computation, they often discard the temporal component of events. Graph-based methods li2021graph ; schaefer2022aegnn ; Zubic_2025_CVPR transform spatio-temporal events into graphs and introduce local update rules for event-by-event processing, thereby capturing both spatial and temporal information. However, graph-based methods are known to suffer from over-smoothing, which limits their scalability and performance on intricate patterns. An attention-based, memory-augmented method was proposed in kamal2023associative to process batched events at each timestep. The LSTM-based Matrix-LSTM cannici2020differentiable method learns and recurrently updates a representation with a pixel-wise LSTM for downstream tasks. FARSE-CNN santambrogio2024farse integrates LSTMs within the neurons of convolutional layers to capture spatio-temporal information and employs local update rules. Our work is most related to the A2S method ALERT-Transformer martin2024alert . In martin2024alert , a patch-wise EventNet is used for asynchronous feature encoding, and the feature is forwarded to a synchronous Transformer for downstream tasks. Our method follows the same A2S paradigm but introduces a new architecture with improved expressivity, solving the problem of limited expressivity brought up by the EventNet encoder.
2.2 Event data-based self-supervised representation learning
@@
Event camera data-based self-supervised representation learning can be broadly categorized into methods that learn only from events and those that incorporate other modalities. For event-only pretraining, MEM klenk2024masked uses the masked image modeling method bao2021beit to event histograms for pretraining. Other works barchid2023exploring ; yang2024event utilize contrastive pretraining with joint embedding architectures chen2020simple ; he2020momentum on event histograms. These methods typically treat event histograms as images and adapt established image pretraining strategies to the event camera domain. When learning with other modalities, event data are often paired with conventional intensity images for representation learning. In yang2023event , contrastive pretraining is explored on event data synthetically converted from RGB images and their source images. EVRepSL qu2024evrepsl learns representations by exploiting the relationship between intensity images and events recorded simultaneously by ATIS cameras. These approaches rely on the correlations between paired visual modalities to guide representation learning.
@@ 3 Method
Our framework’s design leverages the event-language analogy. We first analyze their commonalities and distinctions. On one hand, events and language are similar in their (i) sequential nature and (ii) incremental manner of context contribution. Events convey local intensity changes, similar to language tokens that incrementally build the context (Figure 2). On the other hand, key distinctions exist in their (i) information density and (ii) spatial locality. An individual language token has explicit semantics. In contrast, events only record pixel-level intensity changes and require temporal aggregation to become informative.
Figure 2: Parallels between event data and language: their sequential nature and incremental manner.
As shown in Figure 3, our A2S framework first transforms events into tokens and embeds them with their spatial and temporal attributes. These embeddings are then asynchronously encoded into representations. The representations are learned via self-supervision during training and are generated event-by-event at inference for real-time downstream applications.
Figure 3: Architecture of the A2S framework: (a) Tokenizing and embedding events. (b) Asynchronous encoding via LA. (c) MRP and NRP for self-supervised representation learning.
@@ 3.1 Architecture: asynchronous encoder
Key challenges in designing an asynchronous event encoder encompass: (i) allowing recurrent inference for real-time representation updates and fast parallel training on long event sequences, (ii) continuously encoding events with variable temporal intervals, and (iii) efficiently and expressively aggregating information into the representation. To address the first challenge, drawing from the event-language analogy, we find that recent linear attention (LA) in sequential language modeling supports both recurrent inference and parallel training. Building on LA, we incorporate temporal embedding, MVHS as output, and patch-wise encoding to resolve the other two challenges.
3.1.1 Event tokenization
We define an event sequence as =$\mathcal{E}=\{e_{i}=\left(t_{i},x_{i},y_{i},p_{i}\right),i=1,2,\dots\}$, where $t,x,y,p\in\mathbb{R}$ signify timestamp, coordinates, and polarity of events, with $p\in\{0,1\}$ indicating the two polarities. Each event’s spatial component is tokenized using a bijection $\operatorname{Tok}$ as:
@@ 3.1.1 Event tokenization
where $H,W$ denote the height and width of the spatial area containing the event sequence. The vocabulary size is thus $2\times H\times W$. This mapped token signifies the event’s spatial location.
3.1.2 Embedding events with varied temporal intervals
Each tokenized event $e_{i}=\left(t_{i},\operatorname{Tok}\left(x_{i},y_{i},p_{i}\right)\right)$ is then embedded as an input vector $\bm{x}_{i}\in\mathbb{R}^{D}$ to the encoder. For event tokens that represent spatial locations, we employ a learnable linear embedding on the tokens and get $\operatorname{Emb}_{\text{spatial}}\left(e_{i}\right)\in\mathbb{R}^{D}$. Events are inherently asynchronous, so temporal information must be embedded in the input. Considering that event cameras generate events at high temporal resolution, the absolute timestamp (in $\mu s$) quickly grows large in continuous operation; embedding the absolute timestamp would push the model beyond the range observed during the pretraining period, leading to the same kind of length-extrapolation failures seen in language models press2021train ; su2024roformer ; chen2023extending . We therefore embed the time difference $\Delta t_{i}=t_{i}-t_{i-1}$ instead of the absolute timestamp, using the sinusoidal embedding vaswani2017attention on the $\Delta t_{i}$ as:
$\operatorname{Emb}_{\text{temporal}}\!\left(e_{i}\right)_{k}=\begin{cases}\displaystyle\sin({\Delta t_{i}}/{10000^{\frac{2k}{D}}}),&\text{if }k\text{ even},\\ \displaystyle\cos({\Delta t_{i}}/{10000^{\frac{2k}{D}}}),&\text{if }k\text{ odd},\end{cases}\qquad k=0,1,\dots,D-1.$ (2)
The final embedding is then formed by summing the spatial and temporal embeddings:
$\operatorname{Emb}\left(e_{i}\right)=\operatorname{Emb}_{\text{spatial}}\left(e_{i}\right)+\operatorname{Emb}_{\text{temporal}}\!\left(e_{i}\right)\in\mathbb{R}^{D}.$ (3)
@@ 3.1.3 Modeling event sequence with linear attention
Background: Linear attention in RWKV-6. The asynchronous encoder is based on RWKV-6, a high-performance linear attention architecture for language modelingpeng2024eagle . For an input sequence $\{\bm{x}_{i}\in\mathbb{R}^{D\times 1}\}_{i=1}^{T}$ of length $T$, a single RWKV-6 block performs token mixing (TM) and channel mixing (CM).
In TM, inputs are first mapped to $\{\bm{r}_{i},\bm{w}_{i},\bm{k}_{i},\bm{v}_{i}\in\mathbb{R}^{D\times 1}\}_{i=1}^{T}$. and subsequently mixed by the $\bm{rwkv}$ operator:
@@
where $\bm{u}\in\mathbb{R}^{D\times 1}$ is a learnable parameter. This $\bm{rwkv}$ operator, involving only element-wise prefix sums and products, enables fast training via parallel scan martin2018parallelizing . Equation 4 can be expressed in a recurrent form as:
@@
Consequently, the hidden state $\bm{S}$ can be updated recursively at inference upon the arrival of each new event token, allowing efficient training and recurrent inference on long, continuous event sequences.
@@
The CM layer transforms TM outputs using a gated feed-forward network, similar to that in Transformers vaswani2017attention . Appendix A provides details.
Representing and training with matrix value hidden states (MVHS) as output. Typical NLP tasks use sequence-to-sequence models mapping 1-D token embeddings to 1-D embeddings (or classifying tokens). We propose direct training on the 2-D MVHS, rather than the 1-D output, and use the MVHS as the event representation. Appendix A provides details of the MVHS layer. This practice offers several benefits. First, MVHS, as the hidden state, naturally contains aggregated information. This aggregated representation is precisely what we want to learn, instead of the mapping from event embeddings to event embeddings in standard auto-regressive pretraining (our attempt to predict the next single event failed to converge). Secondly, MVHS yields an expanded memory of size $N\times D_{\text{head}}\times D_{\text{head}}$ without increasing model width $D$, thus improving the representation expressivity and enabling a lightweight architecture for real-time event processing. Using MVHS as output can reduce model size by approximately $D_{\text{model}}/N$ versus 1-D outputs (Appendix F). Moreover, the 2-D structure of MVHS helps to learn fine-grained spatial representations, improving performance over prior 1-D features sekikawa2019eventnet ; martin2024alert .
3.1.4 Patch-wise representation encoding
To exploit the locality of the event, we follow the practice in martin2018parallelizing and use patch-wise representation encoding. For an event camera resolution $\left(H_{\text{sensor}},W_{\text{sensor}}\right)$ and patch size $P$, we divide events into ${H_{\text{sensor}}\times W_{\text{sensor}}}/{P^{2}}$ sequences by coordinates. As Figure 4 illustrates, we learn and encode representations for each patch separately. With patch-wise encoding, the model size of the encoder can be reduced by a factor of approximately #$\#\;patches$, which helps real-time inference (see Appendix F for discussion). Furthermore, as the asynchronous encoder trains on fixed-size $P$ patches, this patch-wise encoding allows application to event cameras with differing resolutions Zubic_2024_CVPR .
3.2 Self-supervised representation learning on asynchronous events
Figure 4: Patch-wise representation encoding. Events are partitioned and encoded by patch. Patch representations could be concatenated for downstream tasks.
The available A2S method learns the representation in a fully supervised manner martin2024alert , leading to limited generalizability to diverse downstream tasks. Here, we propose a self-supervised representation learning method composed of multi-representation prediction (MRP) and next representation prediction (NRP).

### v2

@@ Abstract
Event cameras deliver visual data with high temporal resolution, low latency, and minimal redundancy, yet their asynchronous, sparse sequential nature challenges standard tensor-based machine learning (ML). While the recent asynchronous-to-synchronous (A2S) paradigm aims to bridge this gap by asynchronously encoding events into learned features for ML pipelines, existing A2S approaches often sacrifice expressivity and generalizability compared to dense, synchronous methods. This paper introduces EVA (EVent Asynchronous feature learning), a novel A2S framework to generate highly expressive and generalizable event-by-event features. Inspired by the analogy between events and language, EVA uniquely adapts advances from language modeling in linear attention and self-supervised learning for its construction. In demonstration, EVA outperforms prior A2S methods on recognition tasks (DVS128-Gesture and N-Cars), and represents the first A2S framework to successfully master demanding detection tasks, achieving a 0.477 mAP on the Gen1 dataset. These results underscore EVA’s potential for advancing real-time event-based vision applications.
@@ 1 Introduction
[…]
However, existing A2S work is limited in model expressivity and feature generalizability. On one hand, the asynchronous encoder relies on preliminary models as a compromise for computational efficiency considerations, which demands parallel training and equivalent recurrent inference. Therefore, the A2S method achieves suboptimal performance in complex tasks compared to dense approaches (Gehrig and Scaramuzza, 2023; Gehrig et al., 2019; Peng et al., 2023b; Fan et al., 2025; Zubic et al., 2024; Zubic and Scaramuzza, 2025) with event images. On the other hand, the features are end-to-end learned in a supervised manner and thus are task-specific. This limits their generalizability across diverse downstream applications.
@@
Figure 1: Overview of the proposed EVA A2S framework. It has an asynchronous LA-based encoder (top middle) producing event-by-event updated features for synchronous downstream tasks (right). These features are learned by self-supervision (bottom middle) and are event-by-event updated during inference to be sampled on demand by downstream tasks.
@@
To design an A2S framework with more expressivity and generalizability, we take inspiration from the analogy between events and language (Soydan et al., 2024) and leverage recent advances in natural language processing (NLP). First, events and language are both organized in sequence. Secondly, events record incremental visual changes, similar to words in a sentence that incrementally contribute to the semantic meaning. These two similarities inspire us to treat each event as one token and leverage recent linear attention (LA) architectures for efficient sequence modeling (Katharopoulos et al., 2020; Yang et al., 2023a; Peng et al., 2023a; Sun et al., 2023) and auto-regressive prediction (Radford et al., 2018) for feature learning.
Moreover, we also consider the distinctions between events and language, primarily in information density and spatial locality. One token in NLP has significant semantic meaning, whereas an individual event token offers limited information, gaining significance only when aggregated over time. This steers our focus towards the aggregated global information rather than isolated event instances. Additionally, events typically signify spatially localized changes, contrasting with language, which lacks inherent spatial attributes. This motivates modeling correlations primarily between spatially adjacent events to reduce complexity.
@@
In this work, we introduce a new A2S framework with a more expressive asynchronous encoder, which achieves event-by-event feature updating, and a self-supervised method for generalizable feature learning (Figure 1). The asynchronous encoder is built on the basis of LA, which enables recurrent inference and parallel training. We propose to use the matrix-value hidden state (MVHS) introduced in LA as output to expand the model’s expressivity on the aggregated global information and propose patch-wise encoding to leverage the locality. For ease of implementation and better performance, our model is built on an open source high-performance LA architecture RWKV-6 (Peng et al., 2024), which implements MVHS and is stable to train.
We propose a self-supervised learning (SSL) method to learn generalizable features, which consists of two tasks: multi-representation prediction (MRP) and next-representation prediction (NRP). The former encourages the feature to learn diverse information from multiple handcrafted (converted) representations (Lagorce et al., 2016; Sironi et al., 2018) comprehensively, and the latter, motivated by next-token prediction (NTP) in NLP (Radford et al., 2018), pushes the model beyond simple memorization. Our framework is superior to the previous A2S work (Martin-Turrero et al., 2024) on recognition tasks, and it is the first A2S framework, to the best of our knowledge, that successfully masters demanding detection tasks, achieving a 0.477 mAP on the Gen1 dataset. We name our framework as EVent Asynchronous feature learning (EVA). Our contributions include:
@@
1. An asynchronous LA-based encoder architecture derived from RWKV-6, enabling efficient event-by-event feature updating with improved expressivity.
@@
2. A novel multi-task SSL method learning features generalizable to various downstream tasks.
@@
3. We present EVA, a new A2S framework that combines the asynchronous encoder with self-supervised feature learning, outperforming the previous A2S work on recognition tasks and, for the first time, successfully masters demanding detection tasks with a 0.477 mAP on the Gen1 dataset.
@@ 2.1 Asynchronous event processing
To leverage the sparsity and high temporal resolution of event data, several recent studies have focused on asynchronous event processing. These include convolutional methods (Messikommer et al., 2020; Cannici et al., 2019), graph-based models (Li et al., 2021; Schaefer et al., 2022; Dalgaty et al., 2023; Dampfhoffer et al., 2025; Gehrig and Scaramuzza, 2024; Li et al., 2025), attention-based mechanisms (Kamal et al., 2023), LSTM-based solutions (Cannici et al., 2020; Santambrogio et al., 2024), and point cloud approaches (Sekikawa et al., 2019; Martin-Turrero et al., 2024). Convolutional methods derive local activation update rules in standard (Cannici et al., 2019) or sparse (Messikommer et al., 2020) convolution layers. While they can leverage sparsity and reduce computation, they discard the temporal information. Graph-based methods transform events into spatial-temporal graphs to capture both spatial and temporal information and leverage sparsity. However, graph-based methods show limitations in temporal accumulation (Dampfhoffer et al., 2025). An attention-based, memory-augmented method was proposed in Kamal et al. (2023) to process batched events at each timestep. The LSTM-based Matrix-LSTM (Cannici et al., 2020) method learns and recurrently updates a representation with a pixel-wise LSTM for downstream tasks. FARSE-CNN (Santambrogio et al., 2024) integrates LSTMs within the neurons of convolutional layers to capture spatio-temporal information and employs local update rules. Our work is most related to the A2S method point cloud-based ALERT-Transformer (Martin-Turrero et al., 2024), which uses EventNet (Sekikawa et al., 2019) for asynchronous feature encoding. However, EventNet does not perform hierarchical learning and its expressivity is therefore limited. Our method follows the same A2S paradigm but improves expressivity with a new architecture.
2.2 Language-like event processing
Recently, several works studied processing raw events in a token-wise way. Jiang et al. (2022) processes raw event sequence with Transformer directly, but is limited to the quadratic complexity of softmax attention on long event sequences. To solve this, Schöne et al. (2024) and Soydan et al. (2024) introduce state space models (SSMs) with linear complexity for asynchronous event processing. However, these methods do not discuss the similarity between events and language. Concurrent work (Fang and Panda, 2025) also models raw events with linear attention (Peng et al., 2023a). However, our work focuses on the improvement of feature expressivity under the A2S paradigm, while Fang and Panda (2025) focuses on learning event representations like Word2Vec (Church, 2017).
2.3 Self-supervised learning on event data
@@
Event data SSL can be broadly categorized into methods that learn from events only and those that incorporate other modalities. For event-only SSL, MEM (Klenk et al., 2024) uses the masked image modeling method (Bao et al., 2021) to event images for pretraining. Other works (Barchid et al., 2023; Yang et al., 2024) utilize contrastive learning with joint embedding architectures (Chen et al., 2020; He et al., 2020) on event images. These methods typically convert events to images and adapt established image SSL methods to the event domain. For SSL with other modalities, events are often paired with intensity images. In Yang et al. (2023b), contrastive pretraining is explored on event data synthetically converted from RGB images and their source images. EVRepSL learns representations by exploiting the relationship between paired images and events (Qu et al., 2024).
@@ 3 Method
Our framework leverages the event-language analogy, and we first analyze their similarities and distinctions. On one hand, events and language are similar in their (i) sequential nature and (ii) incremental manner. Events convey local intensity changes, similar to language tokens that incrementally build the context (Figure 2). On the other hand, key distinctions exist in their (i) information density and (ii) spatial locality. An individual language token has explicit semantics. In contrast, events only record pixel-level intensity changes and require temporal aggregation to become informative.
Figure 2: Parallels between events and language: their sequential nature and incremental manner.
As illustrated in Figure 3, our A2S framework first transforms raw events into tokens and embeds them with their spatial and temporal attributes. These embeddings are then asynchronously encoded into features by the RWKV-6-based encoder. The features are learned via SSL during training and are generated event-by-event at inference for real-time downstream applications.
Figure 3: Architecture of the A2S framework: (a) Tokenizing and embedding events. (b) Asynchronous encoding via LA. (c) MRP and NRP for self-supervised feature learning.
@@ 3.1 Architecture: asynchronous encoder
Key challenges in designing an asynchronous event encoder include: (i) allowing recurrent inference for real-time feature updates and parallel training on long event sequences, (ii) continuously encoding events with variable temporal intervals, and (iii) efficiently and expressively aggregating information into the feature. To address the first challenge, drawing from the event-language analogy, we find that recent linear attention (LA) supports both recurrent inference and parallel training. Building on LA, we incorporate temporal embedding, MVHS as output, and patch-wise encoding for the other challenges.
3.1.1 Event tokenization and embedding
We define an event sequence as $\mathcal{E}=\{e_{i}=\left(t_{i},x_{i},y_{i},p_{i}\right),i=1,2,\dots\}$, where $t,x,y,p\in\mathbb{R}$ are the timestamp, coordinates, and polarity of events, with $p\in\{0,1\}$ indicating the two polarities. The spatial component of each event is tokenized using a bijection $\operatorname{Tok}$ as:
@@ 3.1.1 Event tokenization and embedding
where $H,W$ denote the height and width of the spatial area containing the event sequence. The vocabulary size is thus $2\times H\times W$. This mapped token represents the spatial location of the event.
Each tokenized event $e_{i}=\left(t_{i},\operatorname{Tok}\left(x_{i},y_{i},p_{i}\right)\right)$ is then embedded as an input vector $\bm{x}_{i}\in\mathbb{R}^{D}$ to the encoder. For event tokens that represent spatial locations, we employ a learnable embedding layer on the tokens and get $\operatorname{Emb}_{\text{spatial}}\left(e_{i}\right)\in\mathbb{R}^{D}$.
Events are inherently asynchronous, so temporal information must be embedded in the input. Considering that event cameras generate events at high temporal resolution, the absolute timestamp (in $\mu s$) quickly grows large in continuous operation; embedding the absolute timestamp would push the model beyond the range observed during the pretraining period, leading to the same kind of length-extrapolation failures in language models (Press et al., 2021; Su et al., 2024; Chen et al., 2023). We therefore embed the time difference $\Delta t_{i}=t_{i}-t_{i-1}$ instead of the absolute timestamp, and use sinusoidal embedding (Vaswani et al., 2017). The final embedding is then formed by summing the spatial and temporal embeddings.
@@ 3.1.2 Modeling event sequence with linear attention
Background: Linear attention in RWKV-6. The asynchronous encoder is built on RWKV-6, a high-performance linear attention architecture for language modeling(Peng et al., 2024). For an input sequence $\{\bm{x}_{i}\in\mathbb{R}^{D\times 1}\}_{i=1}^{T}$ of length $T$, a single RWKV-6 block performs token mixing (TM) and channel mixing (CM).
In TM, inputs are first mapped to $\{\bm{r}_{i},\bm{w}_{i},\bm{k}_{i},\bm{v}_{i}\in\mathbb{R}^{D\times 1}\}_{i=1}^{T}$ and subsequently mixed by the $\bm{rwkv}$ operator:
@@
where $\bm{u}\in\mathbb{R}^{D\times 1}$ is a learnable parameter. This $\bm{rwkv}$ operator involves only element-wise prefix summations and products and therefore enables parallel training with scan (Martin and Cundy, 2018). Equation 2 can be expressed in a recurrent form:
@@
Consequently, the hidden state $\bm{S}$ can be updated recurrently at inference upon the arrival of each new event, allowing efficient training and recurrent inference on continuous long event sequences.
@@
The CM layer transforms TM outputs using a feed-forward network, similar to Transformers (Vaswani et al., 2017). Appendix B provides details.
Matrix-value hidden states (MVHS) as output. NLP tasks typically involve sequence-to-sequence models mapping 1-D embeddings to 1-D embeddings. We propose to train directly on the 2-D MVHS $\bm{S}$, rather than the 1-D output $\bm{y}$, and use $\bm{S}$ as the event feature, which offers several benefits. First, MVHS, as the hidden state, naturally contains aggregated information. This aggregated feature is precisely what we want, instead of the mapping from embeddings to embeddings in standard auto-regressive pretraining (our attempt to predict the next single event embedding failed to converge). Secondly, MVHS provides an expanded memory of size $N\times D_{\text{head}}\times D_{\text{head}}$ without increasing model width $D$, thus improving the expressivity of the learned feature and enabling a lightweight architecture for real-time event processing.
Figure 4: Patch-wise feature encoding. Events are partitioned and encoded by patch. The patch feature could be concatenated for downstream tasks.
Using MVHS as output can reduce model size by approximately $D_{\text{model}}/N$ versus 1-D outputs (Appendix G). Moreover, the 2-D structure of MVHS helps to learn fine-grained spatial features, improving performance over prior 1-D features (Sekikawa et al., 2019; Martin-Turrero et al., 2024).
3.1.3 Patch-wise encoding of features
To exploit the locality of the event, we follow the practice in Martin and Cundy (2018) and use Patch-wise encoding (PWE) for the feature. Specifically, for an event camera with resolution $\left(H_{\text{sensor}},W_{\text{sensor}}\right)$ and patch size $P$, we separate events to ${H_{\text{sensor}}\times W_{\text{sensor}}}/{P^{2}}$ sequences by their coordinates. As Figure 4 illustrates, we encode a feature for each patch separately. With PWE, the model size of the encoder can be reduced by a factor of approximately the number of patches (#patches\#\;patches), which helps real-time inference (see Appendix G for discussion). Furthermore, since the asynchronous encoder trains on fixed-size patches, this patch-wise encoding allows the application to event cameras with different resolutions.
3.2 Self-supervised feature learning on asynchronous events
We employ SSL to train the asynchronous encoder. Since this training procedure is task-irrelevant, the resulting event feature is not tailored to a certain task, thereby exhibiting better generalizability. The proposed SSL method consists of two tasks: multi-representation prediction (MRP) and next representation prediction (NRP).